# Comparison of Offset and Ratio Weighted Regressions in Tweedie Models with Application to Mid-Term Cancellations

### v1

@@ 2.4 Log-likelihood Function
The GLM employs the maximum likelihood approach to estimate the parameter vector $\beta$ (Nelder and Wedderburn, 1972). The maximum likelihood approach maximizes the likelihood function, which depends on the parameter vector $\beta$ given the observations $y_{1},\ldots,y_{n}$. For a detailed discussion on the theory of maximum likelihood, see Casella and Berger (2024). In the case of the Tweedie distribution, by utilizing the conditional independence between contracts and Definition 2.1.1, the likelihood function is expressed as follows:
@@
$\nabla(\beta)$ $=\left(\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{0}},\ldots,\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{j}}\right)^{t}$
$=\frac{1}{\phi}\left(\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{t}$
$=\frac{1}{\phi}\bm{X}^{t}\bm{D}\bm{R};$
@@ 2.4.1 Iterative Algorithm
Let $\beta_{(k)}$ denote the estimate of $\beta$ at iteration $k$, The IRLS algorithm is defined as follows:
$\beta_{(k+1)}=\beta_{(k)}+\left(I_{n}\left(\beta_{(k)}\right)\right)^{-1}\nabla\left({\beta_{(k)}}\right);$ (2.6)
@@
$I_{n}(\beta)$ $=-\mathrm{E}\!\left[\text{Hess}\left({\beta}\right)\right]$
@@
$=\frac{1}{\phi}\bm{X}^{t}\bm{D}\bm{X},$
where $\text{Hess}\left({\beta}\right)$ is the Hessian matrix. Referring to equation (2.6), we note that the IRLS algorithm relies on the invertibility of the Fisher Information Matrix. The inverse of the Fisher Information Matrix is also used to calculate the asymptotic covariance matrix of the parameter vector estimator. As demonstrated in this paper, comparing the covariance matrices of the parameter vector estimators is crucial for comparing premium estimators in the offset and ratio approaches. Therefore, analyzing the invertibility of the Fisher Information Matrix is of prime importance.
@@
$\forall A\in\mathcal{P}_{m},\forall X\in\mathcal{M}_{m,q}:X^{t}AX\succ 0$ $\Longleftrightarrow X\leavevmode\nobreak\ \text{has rank}\leavevmode\nobreak\ q.$ (2.9)
@@ 2.4.2 Remarks
$\left(I_{n}(\beta)\right)^{-1}\nabla\left(\beta\right)=\left(X^{t}DX\right)^{-1}X^{t}DR.$ (2.10)
@@ 3.1 Offset Approach
As mentioned in Section 1.1, the statistical inference on the parameter vector $\beta$ in offset approach is performed with the lost costs $Y_{1},\ldots,Y_{n}$. To account for the logarithm of a contract’s risk exposure as the offset variable, we assume that $Y_{i}|\mathbf{X}_{i}$ follows the Tweedie distribution with parameters $(\mu_{i},1,\phi,p)$ where $\mu_{i}$ is defined as $\mu_{i}=t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right)$. We also denote an estimate of $\mu_{i}$ in the offset regression by $\widehat{\mu}^{O}_{i}$, obtained as follows:
$\widehat{\mu}^{O}_{i}=\exp\left(\mathbf{X}^{T}_{i}\widehat{\beta}^{O}+\log(t_{i})\right),$ (3.1)
where $\widehat{\beta}^{O}$, an estimate of $\beta$ in offset approach, is obtained using the IRLS algorithm with the following quantities:
@@
    $\nabla^{O}(\beta)$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{t}$ (3.3)
    $=\frac{1}{\phi}\bm{X}^{t}\bm{D}^{O}\bm{R}^{O};$ (3.4)
@@
    $\bm{D}^{O}=\operatorname{diag}\left(t^{2-p}_{i}\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,...,n},\bm{R}^{O}=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n}.$
@@ 3.2 Ratio Approach
$\zeta_{i}=\exp\left(\mathbf{X}^{T}_{i}\beta\right),$ (3.5)
and $\widehat{\zeta_{i}}$ is an estimate of $\zeta_{i}$, equal to $\widehat{\zeta_{i}}=\exp\left(\mathbf{X}^{T}_{i}\widehat{\beta}^{R}\right)$. The vector $\widehat{\beta}^{R}$, an estimate of $\beta$ in ratio approach, is obtained using the IRLS algorithm with the following quantities:
@@
    $\nabla^{R}(\beta)$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}},\ldots,\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}}x_{i,q}\right)^{t}$ (3.7)
    $=\frac{1}{\phi}\bm{X}^{t}\bm{D}^{R}\bm{R}^{R};$ (3.8)
@@
    $\bm{D}^{R}=\operatorname{diag}\left(t_{i}\exp\left\{(2-p)\mathbf{X}^{T}_{i}\beta\right\}\right)_{i=1,...,n},\bm{R}^{R}=\left(\frac{z_{i}}{\zeta_{i}}-1\right)=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n},$
    where $\mu_{i}=t_{i}\zeta_{i}=t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right).$
@@ 3.3 Link between Offset and Ratio Approaches
    In the offset approach, $Y_{i}|\mathbf{X}^{T}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)$. We can examine the ratio of the loss cost to the risk exposure $t_{i}$, denoted as $Z_{i}=\frac{Y_{i}}{t_{i}}$, and apply the Tweedie Invariance Property. It can then be demonstrated that $Z_{i}|\mathbf{X}^{T}_{i}$ follows a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{2-p},\phi,p)$. More formally, we have the following equivalence:
    $Y|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)\equiv Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{2-p},\phi,p).$
@@
    $\bm{D}=\operatorname{diag}\left(t^{2-p}_{i}\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,\ldots,n}\equiv\bm{D}^{O}$
    This result provides an additional verification that these two regression approaches are equivalent, leading to identical inference outcomes for the parameter vector $\beta$.
@@
    Similarly, in the ratio approach, where $Z_{i}|\mathbf{X}^{T}_{i}$ is assumed to follow a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)$, we can apply the Tweedie Invariance Property to demonstrate that $Y_{i}|\mathbf{X}^{T}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{p-1},\phi,p)$. This equivalence is expressed as:
    $Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)\equiv Y|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{p-1},\phi,p).$
@@
    $\bm{D}=\operatorname{diag}\left(t_{i}\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,\ldots,n}\equiv\bm{D}^{R}.$
@@
$\text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{2-p},\phi,p)\ \ \ \text{{versus}}\ \ \ \text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)$
@@ 3.4 Weights Analysis
$\bm{D}^{R}=\bm{D}^{O}=\operatorname{diag}\left(\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,...,n}.$
@@ 4 Comparison of the Approaches
According to our analysis presented in Section 3.3, for a given value of the variance parameter $p$, the difference between weights $w_{i}^{O}$ and $w_{i}^{R}$, of the offset and ratio approaches respectively, depends on the risk exposure $t_{i}$. This difference is significant for contracts with shorter exposure periods, while it becomes negligible for those with exposure periods close to one year. However, a limitation of this analysis is that it does not conclusively determine which approach may be more appropriate for insurance pricing.
Our strategy to compare approaches involves utilizing the distribution of the maximum likelihood estimators of the parameter vectors in the offset and ratio approaches to compare the premium estimators [^1] We will use the expression premium estimator instead of estimated premium throughout this discussion, as the estimated premium refers to the value of a single observation derived from the premium estimator, which is a more general mathematical expression. in these two regression methods. We begin by comparing the asymptotic expected values and variances of the premium estimators in both regression approaches. This comparison utilizes the positive definite matrix property described in equation (2.8), which enables a direct evaluation of the asymptotic covariance matrices of the parameter vector estimators in the offset and ratio regression approaches. We conclude our analysis by applying the concept of stochastic order (or stochastic dominance), as discussed in Shaked and Shanthikumar (2007).
4.1 Notations and assumptions
We continue to consider $n$ independent contracts, using the same notations and assumptions as in previous sections. To facilitate comparison, we define the response variable consistently across both approaches:
  • Offset approach: $Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\zeta_{i},t^{2-p}_{i},\phi,p)$ for $i=1,\ldots,n$,
  • Ratio approach: $Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\zeta_{i},t_{i},\phi,p)$ for $i=1,\ldots,n$,
where $\zeta_{i}$ is the premium for the $i$-th contract without adjustment for risk exposure and is linked to the score $\mathbf{X}^{T}_{i}\beta$ as specified in Equation (4.1):
$\zeta_{i}=\exp\left(\mathbf{X}^{T}_{i}\beta\right).$ (4.1)
Using this formulation, the premium estimators for $\zeta_{i}$ in the offset and ratio approaches, denoted by $\bm{\widehat{\zeta}^{O}}_{i}$ and $\bm{\widehat{\zeta}^{R}}_{i}$, are obtained as follows:
$\bm{\widehat{\zeta}^{O}}_{i}$ $=\exp\left(\mathbf{X}^{T}_{i}\bm{\widehat{\beta}^{O}}\right);$ (4.2)
$\bm{\widehat{\zeta}^{R}}_{i}$ $=\exp\left(\mathbf{X}^{T}_{i}\bm{\widehat{\beta}^{R}}\right);$ (4.3)
Where $\bm{\widehat{\beta}^{O}}$ and $\bm{\widehat{\beta}^{R}}$ are the estimators of $\beta$ in the offset and ratio approaches, respectively.
4.2 Expected Values and Variances Comparison of Premium Estimators
To compare the offset and ratio regression approaches, a valuable direction is to examine the entire distribution of the premium estimator for each approach. However, comparing the expected values and variances of the premium estimators may be sufficient for this analysis. Indeed, when evaluating competing estimators, the optimal choice is generally the one with the smallest bias and variance. Therefore, our quantities of interest are the expected values and variances of the premium estimators for both approaches, as outlined in the following propositions.

### v2

@@ 2.4 Log-likelihood Function
The GLM employs the maximum likelihood approach to estimate the parameter vector $\bm{\beta}$ (Nelder and Wedderburn, 1972). The maximum likelihood approach maximizes the likelihood function, which depends on the parameter vector $\bm{\beta}$ given the observations $y_{1},\ldots,y_{n}$. For a detailed discussion on the theory of maximum likelihood, see Casella and Berger (2024). In the case of the Tweedie distribution, by utilizing the conditional independence between contracts and Definition 2.1.1, the likelihood function is expressed as follows:
@@
$\nabla(\bm{\beta})$ $=\left(\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{0}},\ldots,\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{j}}\right)^{T}$
$=\frac{1}{\phi}\left(\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{\top}$
$=\frac{1}{\phi}\bm{X}^{\top}\bm{D}\bm{R};$
@@ 2.4.1 Iterative Algorithm
Let $\bm{\beta}_{(k)}$ denote the estimate of $\bm{\beta}$ at iteration $k$, The IRLS algorithm is defined as follows:
$\bm{\beta}_{(k+1)}=\bm{\beta}_{(k)}+\left(I_{n}\left(\bm{\beta}_{(k)}\right)\right)^{-1}\nabla\left({\bm{\beta}_{(k)}}\right);$ (2.6)
@@
$I_{n}(\bm{\beta})$ $=-\mathrm{E}\!\left[\text{Hess}\left({\bm{\beta}}\right)\right]$
@@
$=\frac{1}{\phi}\bm{X}^{\top}\bm{D}\bm{X},$
where $\text{Hess}\left({\bm{\beta}}\right)$ is the Hessian matrix. Referring to equation (2.6), we note that the IRLS algorithm relies on the invertibility of the Fisher Information Matrix. The inverse of the Fisher Information Matrix is also used to calculate the asymptotic covariance matrix of the parameter vector estimator. As demonstrated in this paper, comparing the covariance matrices of the parameter vector estimators is crucial for comparing premium estimators in the offset and ratio approaches. Therefore, analyzing the invertibility of the Fisher Information Matrix is of prime importance.
@@
$\forall A\in\mathcal{P}_{m},\forall X\in\mathcal{M}_{m,q}:X^{\top}AX\succ 0$ $\Longleftrightarrow X\penalty 10000\ \text{has rank}\penalty 10000\ q.$ (2.9)
@@ 2.4.2 Remarks
$\left(I_{n}(\bm{\beta})\right)^{-1}\nabla\left(\bm{\beta}\right)=\left(X^{\top}DX\right)^{-1}X^{\top}DR.$ (2.10)
@@ 3.1 Offset Approach
As mentioned in Section 1.1, the statistical inference on the parameter vector $\bm{\beta}$ in offset approach is performed with the loss costs $Y_{1},\ldots,Y_{n}$. To account for the logarithm of a contract’s risk exposure as the offset variable, we assume that $Y_{i}|\mathbf{X}_{i}$ follows the Tweedie distribution with parameters $(\mu_{i},1,\phi,p)$ where $\mu_{i}$ is defined as $\mu_{i}=t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)$. We also denote an estimate of $\mu_{i}$ in the offset regression by $\widehat{\mu}^{O}_{i}$, obtained as follows:
$\widehat{\mu}^{O}_{i}=\exp\left(\mathbf{X}^{\top}_{i}\widehat{\bm{\beta}}^{O}+\log(t_{i})\right),$ (3.1)
where $\widehat{\bm{\beta}}^{O}$, an estimate of $\bm{\beta}$ in offset approach, is obtained using the IRLS algorithm with the following quantities:
@@
    $\nabla^{O}(\bm{\beta})$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{\top}$ (3.3)
    $=\frac{1}{\phi}\bm{X}^{\top}\bm{D}^{O}\bm{R}^{O};$ (3.4)
@@
    $\bm{D}^{O}=\operatorname{diag}\left(t^{2-p}_{i}\exp\left((2-p)\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)_{i=1,...,n},\bm{R}^{O}=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n}.$
@@ 3.2 Ratio Approach
$\zeta_{i}=\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),$ (3.5)
and $\widehat{\zeta_{i}}$ is an estimate of $\zeta_{i}$, equal to $\widehat{\zeta_{i}}=\exp\left(\mathbf{X}^{\top}_{i}\widehat{\bm{\beta}}^{R}\right)$. The vector $\widehat{\bm{\beta}}^{R}$, an estimate of $\bm{\beta}$ in ratio approach, is obtained using the IRLS algorithm with the following quantities:
@@
    $\nabla^{R}(\bm{\beta})$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}},\ldots,\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}}x_{i,q}\right)^{\top}$ (3.7)
    $=\frac{1}{\phi}\bm{X}^{\top}\bm{D}^{R}\bm{R}^{R};$ (3.8)
@@
    $\bm{D}^{R}=\operatorname{diag}\left(t_{i}\exp\left\{(2-p)\mathbf{X}^{\top}_{i}\bm{\beta}\right\}\right)_{i=1,...,n},\bm{R}^{R}=\left(\frac{z_{i}}{\zeta_{i}}-1\right)=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n},$
    where $\mu_{i}=t_{i}\zeta_{i}=t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right).$
@@ 3.3 Link between Offset and Ratio Approaches
    In the offset approach, $Y_{i}|\mathbf{X}^{\top}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),1,\phi,p)$. We can examine the ratio of the loss cost to the risk exposure $t_{i}$, denoted as $Z_{i}=\frac{Y_{i}}{t_{i}}$, and apply the Tweedie Invariance Property. It can then be demonstrated that $Z_{i}|\mathbf{X}^{\top}_{i}$ follows a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{2-p},\phi,p)$. More formally, we have the following equivalence:
    $Y|\mathbf{X}^{\top}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),1,\phi,p)\equiv Z|\mathbf{X}^{\top}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{2-p},\phi,p).$
@@
    D = diag( t^2 - p_i exp( (2 - p) X^⊤_i β ) )_i= 1,…,n ≡D^O
    This result provides an additional verification that these two regression approaches are equivalent, leading to identical inference outcomes for the parameter vector $\bm{\beta}$.
@@
    Similarly, in the ratio approach, where $Z_{i}|\mathbf{X}^{\top}_{i}$ is assumed to follow a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i},\phi,p)$, we can apply the Tweedie Invariance Property to demonstrate that $Y_{i}|\mathbf{X}^{\top}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{p-1},\phi,p)$. This equivalence is expressed as:
    $Z|\mathbf{X}^{\top}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i},\phi,p)\equiv Y|\mathbf{X}^{\top}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{p-1},\phi,p).$
@@
    $\bm{D}=\operatorname{diag}\left(t_{i}\exp\left((2-p)\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)_{i=1,\ldots,n}\equiv\bm{D}^{R}.$
@@
$\text{Tweedie}(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{2-p},\phi,p)\ \ \ \text{{versus}}\ \ \ \text{Tweedie}(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i},\phi,p)$
@@ 3.4 Weights Analysis
$\bm{D}^{R}=\bm{D}^{O}=\operatorname{diag}\left(\exp\left((2-p)\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)_{i=1,...,n}.$
@@ 4 Comparison of the Approaches
According to the analysis presented in Section 3.3, for a fixed value of the variance parameter $p$, the difference between the weights $w_{i}^{O}$ and $w_{i}^{R}$ associated with the offset and ratio approaches, respectively, depends on the risk exposure $t_{i}$. This difference is substantial for contracts with short exposure periods, while it becomes negligible for contracts with exposure periods close to one year. However, a limitation of this analysis is that it does not, by itself, allow us to determine which approach is more appropriate for insurance pricing.
To address this issue, we proceed in two steps. First, we compare the parameter vector estimators derived under the two approaches, relying on the quasi-maximum likelihood framework of White (1982). Second, using a numerical illustration, we highlight the practical merits of the ratio approach relative to the offset approach in the context of automobile insurance pricing. To this end, we consider as the response variable the loss cost $Y$ per unit of risk exposure $t$, namely $Z=\frac{Y}{t}$, and assume that the true data-generating process of $Z$ is characterized by a density function $g$.
For all $\bm{x}\in\mathbb{R}^{q+1}$ and $t\in(0,1]$, representing the covariate vector (risk characteristics) and the risk exposure, respectively, the true expected value of $Z$ is given by
$\zeta^{\text{True}}=\int_{0}^{+\infty}z\,g\!\left(z\mid\bm{\beta}^{\text{True}},t\right)\,dz=\exp\!\left\{\bm{x}^{\top}\bm{\beta}^{\text{True}}\right\}<+\infty,$
where $\bm{\beta}^{\text{True}}\in\mathbb{R}^{q+1}$ denotes the true parameter vector.To clarify our strategy, we also recall the definition of estimator consistency, following Casella and Berger (2024). An estimator $\widehat{\bm{\beta}}$ of $\bm{\beta}^{\text{True}}$ is said to be consistent (or convergent in probability) if
$\widehat{\bm{\beta}}\xrightarrow[]{\mathbb{P}}\bm{\beta}^{\text{True}}\quad\text{as }n\to+\infty,$
that is, if the following condition holds:
$\lim_{n\to\infty}\Pr\!\left(\|\widehat{\bm{\beta}}-\bm{\beta}^{\text{True}}\|>\epsilon\right)=0,\quad\text{for all }\epsilon>0.$
4.1 Parameter Vector Estimators Comparison
Recall that the estimation of the true parameter vector $\bm{\beta}^{\text{True}}$ under the two approaches is carried out by assuming a Tweedie distribution with different weights in each case: the offset approach uses $w=t^{2-p}$, whereas the ratio approach uses $w=t$. We denote the corresponding density by $f$, which is defined as
$f(z\mid\bm{\beta},t)=\exp\!\left(\frac{w}{\phi}\left(\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}\}}{1-p}\,z-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}\right)+a(z,w,\phi)\right),\qquad z\geq 0,\qquad\bm{\beta}\in\mathbb{R}^{q+1}.$
We then define the Kullback–Leibler (KL) divergence between the true density $g(\cdot)$ and the density $f(\cdot)$ as
$KL(\bm{\beta})$ $=$ $\mathbb{E}\!\left[\log\frac{g(z\mid\bm{\beta}^{\text{True}},t)}{f(z\mid\bm{\beta},t)}\right]$
$=$ $\int_{0}^{+\infty}\log\!\big(g(z\mid\bm{\beta}^{\text{True}},t)\big)\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz$
$\quad-\int_{0}^{+\infty}\log\!\big(f(z\mid\bm{\beta},t)\big)\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz.$